If Veronica Vaughn spends all of her daily income on cigarettes and Yoo Hoo, she can afford 10 packs of cigarettes and 10 bottles of Yoo Hoo. She can also afford 6 packs of cigarettes and 22 bottles of Yoo Hoo.

(a)  What is the relative price of cigarettes in terms of bottles of Yoo-Hoo (i.e, what is the ratio of prices, Pc/Py?)

(b)  Exactly how much income does Veronica earn in one week if the price of cigarettes is $6? Write a budget equation for Veronica that is a function of the pack of cigarettes, C, and the number of bottles of Yoo Hoo, Y.

(c)  Draw veronica’s daily budget set with packs of cigarettes on the x-axis and bottles of Yoo Hoo on the y-axis.


1
Expert's answer
2021-11-11T13:27:00-0500

Solution:

a.). Budget constraint: PcC + PyY = M

M = Income

10Pc + 10Py = M

6Pc + 22Py = M

4Pc – 2Py = M

4Pc = 2Py

2Pc = Py

Py = 2Pc

Price ratio = "\frac{Pc}{Py} = \frac{Pc}{2Pc} = 2"

 

b.). PcC + PyY = M

Pc = 6

60 + 10Py = M

36 + 22Py = M

24 – 12Py = 0

24 = 12Py

Py = 2

M = PcC + PyY

M = 6(10) + 2(10) = 60 + 20 = 80

Income (M) = 80

 

c.). Veronica’s budget set is drawn as below: